If a and h are real numbers, find the following values for the given function.
f(x) = x2 − x + 5
(a)
f(a)
(b)
f(−a)

(c)
−f(a)

(d)
f(a + h)


(e)
f(a) + f(h)
(f)
f(a + h) − f(a)
h
, h ≠ 0

Answer by mananth(16946)

If a and h are real numbers, find the following values for the given function.
f(x) = x2 − x + 5
(a) f(a)
f(a)=a ^2−a+5
(b)f(−a)
f(-a) = a^2+a+5
(c) −f(a)
-f(a) = -(a ^2−a+5) = -a^2+a-5
(d) f(a + h)
f(a+h) = (a+h)^2-(a+h) +5
= a^2+2ah+h^2-a-h+5

(e) f(a) + f(h)
f(a)+f(h) = a ^2−a+5 +h^2-h+5
(f) (f(a + h) − f(a))/h h ≠ 0
(f) ((f(a + h) − f(a))/h= (a^2+2ah+h^2-a-h+5 - a ^2+a-5))/h